I'm stuck in the following questions

Please help

14. Use a mm marking on the ruler to estimate the diameter of the field of view at 100x .what is it?

16. Assume that the size of the field of view always changes in proportion to ny change in magnification.

A)how many times difference is there in magnification between  40X and 400 X?

B) what is the diameter of the field of view at 40X?

D) use this information to calculate the approximate diameter diameter of the field of view at 400 X magnification. show how you figured it out.

18.At 100x what happens to the focus of the Lower hair if you keep the upper hair in focus?

Thanks for your help
Answer #1

14.

Diameter of field of view (mm) = field number of eyepiece/magnification of objective lens

Field number is written on the eyepiece. (You need to check, I have given an example).

For field number of 20 mm,

Diameter of field of view = 20/100 = 0.2 mm

16.

A. Divide the two magnifications given,

400x/40x = 10

It means 400x magnification is 10 times greater than 40x.

We can do reverse as well.

40x/400x = 1/10

It means 40x magnification is 10 times smaller than 400x.

B. Diameter of field of view = 20/40 = 0.5 mm

(Using same field number as in 14)

As magnification increases, the diameter of field of view decreases.

D. Diameter (D) of field of view is inversely proportional to magnification (M), if same eyepiece is used.

So, the formula will become,

D1/D2 = M2/M1

D1 = ?, M1 = 400x

D2 = 0.2 (done in last part), M2 = 40x

D1/0.5 = 40x/400x

D1 = 0.05 mm for 400x magnification.

18. The point which is at focus will always appear sharp to us and other unfocussed points will be blurred. So lower hair will appear blur and upper hair will be sharp.

Please rate.